OpenStudy (anonymous):

the perimeter of a rectangle is 70m The length is 8m more than twice the width find the dimension what is length ---m what is width ---m

OpenStudy (anonymous):

width= 9m length= 26m

OpenStudy (anonymous):

could you show me how you arrived at this answer please thanks for your time and help

OpenStudy (anonymous):

let x be the breadth, perimeter of rectangle=2(length+breadth)=2(2x+8+x)=70 6x+16=70 6x=70-16 6x=54 x=54/6=9 length=(9*2)+8=26m breadth=9m

OpenStudy (anonymous):

thanks hakkunamatata you're awesome and prompt